# A Systematic Review of Aspect-based Sentiment Analysis (ABSA): Domains, Methods, and Trends thanks: Preprint manuscript

### v4

@@ 3.3 Note on “domain” mapping
We then manually grouped the research and dataset domains into 19 common categories used for analysis. More details and examples on domain mapping are available in Subsection A.3 in Appendix A.
@@ 4.1 Results for RQ1
To answer RQ1, we examined the distribution of reviewed studies by their research (application) domains, dataset domains, and the relationship between the two. From the 519 reviewed studies, we recorded 218 datasets, 19 domain categories (15 research domains and 17 dataset domains), and obtained 1179 distinct “study - dataset” pairs and 630 unique “study - dataset domain” combinations. The key results are summarised below and presented in Table 2 and Figure 2.
@@
Non-specific 339* 65.32% 98 15.56%
Product/service review 126 24.28% 447* 70.95%
@@ 4.1.1 Detailed Results for RQ1
As to the dataset domains, Table 2 suggests that among the 630 unique “study - dataset domain” pairs, the majority (70.95%, N=447) are in the “product/service review” category, followed by 15.56% (N=98) in “Non-specific”. The third place is shared by two magnitude-smaller categories: “student feedback/ education review” (3.02%, N=19) and “video/movie review” (3.02%, N=19). The numbers of studies with datasets from the Healthcare/medicine (1.43%, N=9) and Politics/policy-reaction (0.79%, N=5) domains were again single-digit. Moreover, nearly half of the unique datasets in the public domains were created by the authors for the first time: 5/9 in Healthcare/medicine, 2/4 in Politics/policy-reaction, and 8/12 in Student feedback/ Education review.
Furthermore, to understand the dataset diversity across samples and domains, we grouped the 1179 unique “study - dataset” pairs by “research-domain, dataset-domain, dataset-name” combinations and zoomed into the 757 entries with ten or more study counts each. As shown in Table 3 and illustrated in Figure 4, among these 757 unique combinations, 95.77% (N=725) are in the “non-specific” research domain, of which 90.48% (N=656) used “product/service review” datasets. Most interestingly, these 757 entries only involve 12 distinct datasets of which 10 were product and service reviews, and 78.20% (N=592) are taken up by the four SemEval datasets from the early pioneer work [31, 32, 33] mentioned in Subsection 1.1: SemEval 2014 Restaurant, SemEval 2014 Laptop (these two alone account for 50.33% of all 757 entries), SemEval 2016 Restaurant, and SemEval 2015 Restaurant. This finding echos [96, 36]: “The SemEval challenge datasets … are the most extensively used corpora for aspect-based sentiment analysis” [36,  p.4]. Meanwhile, the top dataset used under “product/service review” research and dataset domains is the original product review dataset created by the researchers. [36, 97] provides a detailed introduction to the SemEval datasets.
@@ 4.2 Results for RQ2
For RQ2, we examined the 13 recorded subtasks and 807 unique “study - subtask” pairs to identify the most explored ABSA subtasks and subtask combinations across the 519 reviewed studies. As shown in Table 4, 32.37% (N=168) of the studies developed full-ABSA solutions through the combination of AE and ASC, and a similar proportion (30.83%, N=160) focused on ASC alone, usually formulating the research problem as contextualised sentiment analysis with given aspects and the full input text. Only 15.22% (N=79) of the studies solely explored the AE problem. This is consistent with the number of studies by individual subtasks shown in Table 5, where ASC is the most explored subtask, followed by AE and ACD. Moreover, Figure 5 reveals a small but noticeable rise in composite subtasks ASTE (N=1, 5 and 10 in 2017, 2021, 2022) since 2020 and a decline in ASC and AE around the same period.
Table 4: Number of Studies by ABSA Subtask Combinations (N=519)
Subtask Combination Count of Studies % of Studies
AE, ASC 168 32.37%
ASC 160 30.83%
AE 79 15.22%
ACD, ASC 16 3.08%
ASTE 15 2.89%
AE, ACD, ASC 13 2.50%
AE, OE, ASC 9 1.73%
AE, ACD 9 1.73%
AOPE 9 1.73%
ACD 7 1.35%
AE, OE 7 1.35%
AE, OE, ACD 4 0.77%
AE, OE, ACD, ASC 4 0.77%
ASQE 2 0.39%
AOPE, ASC 2 0.39%
OE 1 0.19%
OTE (opinion target extraction), ASC 1 0.19%
Aspect-based embedding, ACD, ASC 1 0.19%
ASC, OE 1 0.19%
Aspect and synthetic sample discrimination 1 0.19%
AE, OE, ASC, AOPE, ASTE 1 0.19%
AOPE, ASC, ACD 1 0.19%
AOPE, ACD 1 0.19%
AE, review-level SA 1 0.19%
ACD, AE, ASC 1 0.19%
AE, ASC, Aspect-based sentence segmentation 1 0.19%
AE, ASC, Aspect-based embedding 1 0.19%
AE, ASC, ACD 1 0.19%
AE, ACD, OE 1 0.19%
Data augmentation 1 0.19%
TOTAL 519 100.00%
Table 5: Number of Studies by Individual ABSA Subtasks (N=807)
Individual Subtask Count of Studies % of Studies
ASC 381 47.21%
AE 300 37.17%
ACD 59 7.31%
OE 28 3.47%
ASTE 16 1.98%
AOPE 14 1.73%
ASQE 2 0.25%
Aspect-based embedding 2 0.25%
Aspect and synthetic sample discrimination 1 0.12%
OTE (opinion target extraction) 1 0.12%
Aspect-based sentence segmentation 1 0.12%
Data augmentation 1 0.12%
Review-level SA 1 0.12%
TOTAL 807 100.00%
@@ 4.3.1 Paradigms
Table 6: Number of Studies by Paradigms (N=519). Note the unsupervised category includes non-machine learning approaches)
@@
We further analysed the approaches under each paradigm and focused on three for more details: deep learning (DL), traditional machine learning (ML), and Linguistic and Statistical Approaches. The results are displayed in Tables 7 to  10 and described below.
Table 7: Number of Studies by Deep Learning (DL) Approaches (N=313)
DL Approach Count of Studies % of Studies
RNN, Attention 83 26.52%
Attention 60 19.17%
RNN 31 9.90%
RNN, CNN, Attention 25 7.99%
CNN 20 6.39%
RNN, CNN 19 6.07%
GNN/GCN, Attention 17 5.43%
CNN, Attention 17 5.43%
RNN, GNN/GCN, Attention 17 5.43%
Other (<5% each) 24 7.67%
TOTAL 313 100.00%
@@ 4.4 Approaches
As shown in Tables 7 and  8, among the 519 reviewed studies, 60.31% (N=313) employed DL approaches, and 30.83% (N=160) are DL-only. The DL-only approach is particularly prominent among fully-supervised (47.15%, N=149) and semi-supervised (31.82%, N=7) studies. Supplementing DL with syntactical features is also the second most popular approach in fully-supervised studies (16.77%, N=53).
  1. 1) DL Approaches
Table 7 suggests that the 313 studies involving DL approaches are dominated by Recurrent Neural Network (RNN)-based solutions (55.91%, N=175), of which nearly half used a combination of RNN and the attention mechanism (26.52%, N=83), followed by attention-only (19.17%, N=60) and RNN-only (9.90%, N=31) models. In addition, convolutional and graph-neural approaches (e.g. Convolutional Neural Networks (CNN), Graph Neural Networks (GNN), Graph Convolutional Networks (GCN)) also play smaller but noticeable roles in DL-based ABSA studies, with the former commonly used as an alternative to the sequence models [20, 98, 99] and the latter for injecting external knowledge and relations to the architecture or better capture local and global context [86, 100, 101].
Figure 6 depicts the trend of the main approaches across the publication years. We excluded the one study pre-published for 2023 to avoid confusing trends. It is clear that DL approaches have risen sharply and taken dominance since 2017, mainly driven by the rapid growth in RNN- and attention-based studies. GNN/GCN-based approaches remain small in number but have noticeable growth since 2020 (N=2, 2, 16, 24 in each of 2019-2022, respectively).
@@
Table 9: Number of Studies by Traditional Machine Learning Approaches (N=283)
Traditional ML Count of Studies % of Studies
Support Vector Machine (SVM) 57 20.14%
Conditional Random Field (CRF) 41 14.49%
Latent Dirichlet allocation (LDA) 36 12.72%
Naïve Bayes (NB) 32 11.31%
Random Forest (RF) 24 8.48%
Decision Tree (DT) 18 6.36%
Logistic Regression (LR) 15 5.30%
K-Nearest Neighbors (KNN) 12 4.24%
Multinomial Naïve Bayes (MNB) 10 3.53%
K-means 5 1.77%
Boosting 5 1.77%
Other (N<5) 28 9.89%
TOTAL 283 100.00%
  1. 2) Traditional ML Approaches
Interestingly, as shown in Figure 6 traditional ML approaches remain a steady force over the decades despite the rapid rise of DL methods. Tables 8 and  9 provide some insight into this: Among the 519 reviewed studies, while 60.31% employed DL approaches as mentioned in the previous sub-section, over half (54.53%, N=283) also included traditional ML approaches, with the top 3 being Support Vector Machine (SVM; 20.14%, N=57), Conditional Random Field (CRF; 14.49%, N=41), and Latent Dirichlet allocation (LDA; 12.72%, N=36). Table 8 suggests that among the major paradigms, traditional ML were often used in combination with DL approaches for fully-supervised studies (7.28%, N=23), and along with linguistic rules, syntactic features, and/or lexicons and ontology in hybrid studies (27.40%, N=20). Across all paradigms, traditional ML-only approaches are relatively rare (max N=7).
  1. 3) Linguistic and Statistical Approaches
While Table 8 illustrates the prevalence of fusing ML approaches with linguistic and statistical features or modules, there were 67 studies (12.91% out of the total 519) on pure linguistic or statistical approaches. As shown in Tables 10, although small in number, these non-ML approaches have persisted over time. The most popular combination (34.33%, N=23) was rules built on syntactic features (e.g. POS tags and dependency parse trees) and used along lexicon resources (e.g. domain-specific aspect lists, SentiWordNet [^3] https://github.com/aesuli/SentiWordNet, MPQA [^4] https://mpqa.cs.pitt.edu/). A typical example is using POS tags and/or lexicon resources to narrow the scope of aspect or opinion term candidates, applying further rules based on POS tags or dependency relations for AE or OE, and/or using lexicon resources for candidate pruning, categorisation, or sentiment labelling. The second top combination (23.88%, N=16) is the above-mentioned one plus ontology (e.g. domain-specific ontology, ConceptNet [^5] https://conceptnet.io/, WordNet [^6] https://wordnet.princeton.edu/) to bring in external knowledge of concepts and relations. Pure statistical methods were relatively rare (5.97%, N=4), and mainly included frequency-based methods such as N-gram and TF-IDF, and other statistical modelling methods that were not commonly seen in the ML field.

### v5

@@ 3.3 Note on “domain” mapping
We then manually grouped the research and dataset domains into 19 common categories used for analysis. More details and examples on domain mapping are available in Appendix B.3.
3.4 Phase-2 targeted review on in-context learning
Additionally, generative “foundation models” [84], defined as models with billions of parameters pre-trained on enormous general-purpose data and adaptable to diverse downstream NLP tasks, have become ubiquitous after our SLR data collection (e.g. ChatGPT [85], released in November 2022). We use the term “foundation models” to distinguish them from the earlier pre-trained Large Language Models (LLMs) such as BERT [50], BART [86], and T5 [87], which have relatively fewer parameters and typically require fine-tuning for task adaptation [73]. These generative foundation models brought a new paradigm of “In-context Learning” (ICL) [51, p. 4], where task adaptation can occur solely via conditioning the model on the text input instructions (“prompts”) with zero (“zero-shot ICL”) or few (“few-shot ICL”) examples and no model parameter changes [51, 88]. To capture and analyse this new development while balancing feasibility and currency, we conducted a Phase-2 targeted review in July 2024.
This Phase-2 targeted review focuses solely on the ICL implementations of pre-trained generative models for ABSA tasks, excluding those involving fine-tuning to draw a distinction from other non-ICL deep-learning approaches covered in the SLR. To extend the SLR sample beyond the original extraction time, we conducted a new database search [^2] This new database search followed the same procedures and criteria as the SLR, except that we aborted the SpringerLink search due to persistent database interface search result navigation issues during our data collection period. in July 2024 for studies published from 2022 onwards and removed the ones already included in the SLR sample. The new search results were screened using the SLR criteria described in Table 1 and then combined with the 519 SLR final samples. We then applied an additional filtering condition “Gen-LLM” to all the in-scope ABSA primary studies, which further selected publications with at least one occurrence of any of the following keywords outside the Reference section: “generative”, “in-context”, “in context learning”, “genai”, “bart”, “t5”, “flan-t5”, “gpt”, “chatgpt”, “llama”, and “mistral”. With the help of our automatic screening pipeline detailed in Appendix B.2, we were able to efficiently auto-screen the new search results and re-screen the previous SLR sample for the "Gen-LLM" keywords in less than one hour.
In total, the new search yielded 271 additional in-scope ABSA primary studies from 4359 search results. After applying the “Gen-LLM” filtering condition to the combined 790 in-scope ABSA primary studies, we obtained 208 Phase-2 samples for manual review, which comprised 91 studies from the new search and 117 from the earlier SLR sample, ranging from 2008 to 2024. The Phase-2 targeted review results are presented in Section 4.5. Unless specified otherwise, the results below only refer to those of the SLR.
@@ 4.1 Results for RQ1
To answer RQ1, we examined the distribution of reviewed studies by their research (application) domains, dataset domains, and the relationship between the two. From the 519 reviewed studies, we recorded 218 datasets, 19 domain categories (15 research domains and 17 dataset domains), and obtained 1179 distinct “study-dataset” pairs and 630 unique “study & dataset-domain” combinations. The key results are summarised below and presented in Table 2 and Figure 2. We also list the datasets used by more than one reviewed study in the Appendix Table LABEL:tab:c5_dataset.
@@
Non-specific 339 65.32% 98 15.56%
Product/service review 126 24.28% 447 70.95%
@@ 4.1.1 Detailed Results for RQ1
As to the dataset domains, Table 2 suggests that among the 630 unique “study & dataset-domain” pairs, the majority (70.95%, N=447) are in the “product/service review” category, followed by 15.56% (N=98) in “Non-specific”. The third place is shared by two magnitude-smaller categories: “student feedback/ education review” (3.02%, N=19) and “video/movie review” (3.02%, N=19). The numbers of studies with datasets from the Healthcare/medicine (1.43%, N=9) and Politics/policy-reaction (0.79%, N=5) domains were again single-digit. Moreover, nearly half of the unique datasets in the public domains were created by the authors for the first time: 5/9 in Healthcare/medicine, 2/4 in Politics/policy-reaction, and 8/12 in Student feedback/ Education review.
Furthermore, to understand the dataset diversity across samples and domains, we grouped the 1179 unique “study-dataset” pairs by “research-domain, dataset-domain, dataset-name” combinations and zoomed into the 757 entries with ten or more study counts each. As shown in Table 3 and illustrated in Figure 4, among these 757 unique combinations, 95.77% (N=725) are in the “non-specific” research domain, of which 90.48% (N=656) used “product/service review” datasets. Most interestingly, these 757 entries only involve 12 distinct datasets of which 10 were product and service reviews, and 78.20% (N=592) are taken up by the four SemEval datasets from the early pioneer work [61, 62, 63] mentioned in Subsection 2.4: SemEval 2014 Restaurant, SemEval 2014 Laptop (these two alone account for 50.33% of all 757 entries), SemEval 2016 Restaurant, and SemEval 2015 Restaurant. This finding echos [89, 47]: “The SemEval challenge datasets … are the most extensively used corpora for aspect-based sentiment analysis” [47,  p.4]. Meanwhile, the top dataset used under “product/service review” research and dataset domains is the original product review dataset created by the researchers. [47, 90] provides a detailed introduction to the SemEval datasets.
@@ 4.2 Results for RQ2
For RQ2, we examined the 13 recorded subtasks and 805 unique “study-subtask” pairs to identify the most explored ABSA subtasks and subtask combinations across the 519 reviewed studies. As shown in Figure 5(a), 32.37% (N=168) of the studies developed full-ABSA solutions through the combination of AE and ASC, and a similar proportion (30.83%, N=160) focused on ASC alone, usually formulating the research problem as contextualised sentiment analysis with given aspects and the full input text. Only 15.22% (N=79) of the studies solely explored the AE problem. This is consistent with the number of studies by individual subtasks shown in Figure 5(b), where ASC is the most explored subtask, followed by AE and ACD.
Moreover, Figure 6 reveals a small but noticeable rise in composite subtask ASTE since 2020 (N=1, 5 and 10 in 2017, 2021, 2022) and a decline in ASC and AE around the same period. This could signify a problem formulation shift driven by deep-learning, especially multi-task learning methods for E2E ABSA. Our Phase-2 targeted review findings in Subsection 4.5 add more insights into this.
(a) Number of Studies per ABSA Subtask Combinations (N=519)
(b) Number of Studies per Individual ABSA Subtask (N=805)
Figure 5: Number of Studies by ABSA Subtask
@@ 4.3.1 Paradigms
Table 4: Number of Studies by Paradigm (N=519).
Note: The unsupervised category includes non-ML approaches
@@
We further analysed the approaches under each paradigm and focused on three for more details: deep learning (DL), traditional machine learning (ML), and Linguistic and Statistical Approaches. The results are detailed below and presented in Figure 7 and Tables 5,  6.
(a) Number of Studies by DL Approaches (N=313)
(b) Number of Studies by Traditional ML Approaches (N=283)
Figure 7: Number of Studies using DL and Traditional ML approaches
@@ 4.4 Approaches
As shown in Figure 7(a) and Table 5, among the 519 reviewed studies, 60.31% (N=313) employed DL approaches, and 30.83% (N=160) are DL-only. The DL-only approach is particularly prominent among fully-supervised (47.15%, N=149) and semi-supervised (31.82%, N=7) studies. Supplementing DL with syntactical features is also the second most popular approach in fully-supervised studies (16.77%, N=53).
  • 1) DL Approaches
Figure 7(a) suggests that the 313 studies involving DL approaches are dominated by Recurrent Neural Network (RNN)-based solutions (55.91%, N=175), of which nearly half used a combination of RNN and the attention mechanism (26.52%, N=83), followed by attention-only (19.17%, N=60) and RNN-only (9.90%, N=31) models. The RNN family mainly consists of Long Short-Term Memory (LSTM), Bidirectional LSTM (BiLSTM), and Gated Recurrent Unit (GRU). These neural-networks are featured by sequential processing that captures temporal dependencies of text tokens, and can thus incorporate surrounding text as context for prediction [20, 54]. On the other hand, the sequential nature poses challenges with parallelisation and the exploding and vanishing gradient problems associated with long sequences [74, 20]. Although LSTM and GRU can mitigate these issues somewhat through cell state and memory controls, efficiency and long-dependency challenges still hinder their performance [74, 20, 54]. The attention mechanism complements RNNs by dynamically updating weights across the input sequence based on each element’s relevance to the current task, and thus guides the model to focus on the most relevant elements [74].
In addition, convolutional and graph-neural approaches (e.g. Convolutional Neural Networks (CNN), Graph Neural Networks (GNN), Graph Convolutional Networks (GCN)) also play smaller but noticeable roles in DL-based ABSA studies. While CNN was commonly used as an alternative to the sequence models such as RNNs [20, 91, 92], the graph-based networks (GNN, GCN) were mainly used to model the non-linear relationships such as external conceptual knowledge (e.g. [11]) and syntactic dependency structures (e.g. [55, 39, 31]) that are not well captured by the sequential networks like RNNs and the flat structure of the attention modules. As a result, they inject richer context into the overall learning process [42, 93, 94].
Figure 8 depicts the trend of the main approaches across the publication years. We excluded the one study pre-published for 2023 to avoid confusing trends. It is clear that DL approaches have risen sharply and taken dominance since 2017, mainly driven by the rapid growth in RNN- and attention-based studies. This coincides with the appearance of the Transformer architecture in 2017 [74] and the resulting pre-trained models such as BERT [50] that were a popular embedding choice to be used alongside RNNs in DL and hybrid approaches (e.g., [95, 43]). GNN/GCN-based approaches remain small in number but have noticeable growth since 2020 (N=2, 2, 16, 24 in each of 2019-2022, respectively), suggesting an increased effort to dynamically integrate relational context into the learning process within the DL framework.
@@
  • 2) Traditional ML Approaches
Interestingly, as shown in Figure 8 traditional ML approaches remain a steady force over the decades despite the rapid rise of DL methods. Table 5 and Figure 7(b) provide some insight into this: Among the 519 reviewed studies, while 60.31% employed DL approaches as mentioned in the previous sub-section, over half (54.53%, N=283) also included traditional ML approaches, with the top 3 being Support Vector Machine (SVM; 20.14%, N=57), Conditional Random Field (CRF; 14.49%, N=41), and Latent Dirichlet allocation (LDA; 12.72%, N=36). Table 5 suggests that among the major paradigms, traditional ML were often used in combination with DL approaches for fully-supervised studies (7.28%, N=23), and along with linguistic rules, syntactic features, and/or lexicons and ontology in hybrid studies (27.40%, N=20). Across all paradigms, traditional ML-only approaches are relatively rare (max N=7).
  • 3) Linguistic and Statistical Approaches
While Table 5 illustrates the prevalence of fusing ML approaches with linguistic and statistical features or modules, there were 67 studies (12.91% out of the total 519) on pure linguistic or statistical approaches. As shown in Tables 6, although small in number, these non-ML approaches have persisted over time. The most popular combination (34.33%, N=23) was rules built on syntactic features (e.g. POS tags and dependency parse trees) and used along lexicon resources (e.g. domain-specific aspect lists, SentiWordNet [^3] https://github.com/aesuli/SentiWordNet, MPQA [^4] https://mpqa.cs.pitt.edu/). A typical example is using POS tags and/or lexicon resources to narrow the scope of aspect or opinion term candidates, applying further rules based on POS tags or dependency relations for AE or OE, and/or using lexicon resources for candidate pruning, categorisation, or sentiment labelling (e.g. [96, 7, 97]). The second top combination (23.88%, N=16) is the above-mentioned one plus ontology (e.g. domain-specific ontology, ConceptNet [^5] https://conceptnet.io/, WordNet [^6] https://wordnet.princeton.edu/) to bring in external knowledge of concepts and relations (e.g. [34, 98]). Pure statistical methods were relatively rare (5.97%, N=4), and mainly included frequency-based methods such as N-gram and TF-IDF, and other statistical modelling methods that were not commonly seen in the ML field.
@@
4.5 ICL and generative approach in ABSA - Phase-2 targeted review results
ICL is a subgenre of the DL approach. However, we discuss the relevant results in this separate subsection due to the Phase-2 review’s more focused sample and finer granularity. Despite the trending popularity of the ICL approach in NLP research and applications since 2022 [88], our results suggest that the ABSA research community is just beginning to explore it with caution. Among the 208 ABSA studies from 2008 to 2024 containing at least one occurrence of the “Gen-LLM” keywords, only five (all published in 2024) applied ICL to both composite and traditional ABSA tasks. All of these studies were exploring the performance of foundation models via ICL against other approaches, rather than focusing on an ICL ABSA solution. Table 7 summarises the models, ABSA tasks, and key findings from these studies. Overall, four of the five studies found that zero-shot and even 5-shot ICL on foundation models (mainly GPTs) could not reach the performance of fine-tuned or fully trained DL models, especially those leveraging pre-trained LLMs to fine-tune a contextual-embedding.
Table 7: Studies with In-context Learning (ICL) Approach on ABSA Tasks (N=5 out of 208 samples from 2008-2024)
Note: The Non-ICL Models were trained or fine-tuned unless specified otherwise
Paper Task Non-ICL Models ICL Models ICL approach Result
Zhou et al. (2024) [99] ASQE RoBERTa ChatGPT 5-shot ICL ChatGPT performed worse than almost all non-ICL methods with about 20-30% lower micro-F1
Liu et al. (2024) [37] ASTE BERT ChatGPT Zero-shot, 5-shot ICL ChatGPT 5-shot ICL performed better than 0-shot ICL, but was still around 20 % lower in F1 score than the main method
Su et al. (2024) [100] ASC T5 Llama2-7b-chat, Llama2-13b-chat, ChatGPT-3.5 N.A. ChatGPT 3.5 performed slightly better than the Llama-2 models across datasets, but was still up to about 20% lower than the main model in accuracy and F1
Amin et al. (2024) [101] AE, ASC, OE LSTM, RoBERTa (not fine-tuned) GPT 3.5-Turbo, GPT-4 Zero-shot On AE and ASC tasks, RoBERTa performed significantly better than the two GPT models with up to about 20% higher accuracy. GPT 3.5-Turbo performed better than GPT-4 on AE and ASC tasks, and outperformed all the other models with the OE task with up to 15% higher accuracy
Mughal et al. (2024) [72] ASC, ACSA LSTM, Flan-T5, DeBERTa, PaLM GPT 3.5-Turbo, PaLM-bison Zero-shot PaLM showed the best overall performance on ASC and ASCA tasks in terms of accuracy and F1, closely followed by fine-tuned DeBERTa. However, PaLM had the lowest accuracy and F1 scores with the multi-aspect multi-sentiment (MAMS) dataset, with 21-48% lower accuracy and F1 than fine-tuned DeBERTa
In addition, we identified an emerging trend by examining the Phase-2 review non-ICL samples: Those employing fine-tuned generative LLMs mostly formulated the ABSA tasks as Sequence-to-Sequence (Seq2Seq) text generation problems, with a particular focus on composite tasks such as ASTE and ASQE. As shown in Table 8, within the 208 samples, a total of 18 studies (all from the new search) published in 2022-2024 applied pre-trained generative LLMs with fine-tuning. The majority of these studies used models based on T5 (N=9) and BART (N=5) with the full Transformer [74] encoder-decoder architecture, followed by encoder-only (N=3, BERT and RoBERTa [102]) and decoder-only (N=1, GPT-2 [103]) models. All but two of these 18 studies were on composite ABSA tasks, mainly ASTE and ASQE. Moreover, two studies ([104, 105]) also leveraged the generative capability of these LLMs to augment training data to enrich the fine-tuned embedding.
Compared with this Seq2Seq generation approach, the common applications of pre-trained LLMs in earlier studies from the main SLR sample often formulate the ABSA task as a classification problem [73]. These studies mostly use encoder-only LLMs for their pre-trained representations to fine-tune a contextual embedding [73], which is then connected to other context-injection or relationship-learning modules and a classifier output layer. For instance, Zhang et al. (2022) [43] employed pre-trained BERT with BiLSTM, a Feed-forward Neural Network (FFNN), and CRF. Li et al. (2021) [95] used pre-trained BERT as an encoder and a decoder featuring a GRU. In contrast, the Seq2Seq generative approach can be illustrated by the signature “Generative Aspect-based Sentiment analysis (GAS)” proposed by Zhang et al. (2021) [106], which leveraged the LLM’s pre-trained and fine-tuned encoder module for context-aware embedding and used the fine-tuned decoder module to generate text representations of the label sets (e.g., triplets) or as annotations next to the original input text [106, 73].
[b]
Table 8: Studies with Fine-tuned Generative Large Language Models (LLMs) on ABSA Tasks (N=18 out of 208 samples from 2008-2024)
Paper Task GenAI Model
Hoang et al. (2022) [107] ASQE BART
Kang et al. (2022) [108] OE BART
Gong & Li (2022) [109] ASTE BERT
William & Khodra (2022) [110] ASTE T5
Li et al. (2023) [111] ASQE BART (with BERT for upstream embedding)
Li et al. (2023) [112] ASQE BART
Suchrady & Purwarianti (2023) [113] ASTE, AOPE, AE, OE, UABSA 1 mT5 2
Yu et al. (2023) [114] AESC 3, AOPE, ASTE BART + GAT (Graph Attention Networks)
Yan et al. (2023) [115] ASC,ASTE T5
Yu et al. (2023) [104] AE, OE, ASQE T5 (data generator + self-training) + BERT pair-classifier as descriminator
Lee & Kim (2023) [116] ASTE, ASQE, ASC T5
Liu et al. (2024) [37] ASTE BERT + GPN (Global Pointer Network)
Lee et al. (2024) [117] AE, ASC GPT-2 fine-tuned by LoRA
Dang et al. (2024) [118] AE, OE, ASC, ACD, AOPE, ACSA, ASTE, ACSD, ASQE mT5, ViT5 2
Su et al. (2024) [100] ASC T5
Zhou et al. (2024) [99] ASQE RoBERTa
Wang et al. (2024) [119] AE, OE, ACD, ASC, ACSA, AOPE, ASTE, ASQE, ASPE 4, CSPE 5 T5
Zhang et al. (2024) [105] ASQE T5
  • 1
    UABSA: Unified Aspect-based Sentiment Analysis.
  • 2
    mT5, ViT5: multi-lingual variants of T5.
  • 3
    AESC: Aspect Term Extraction and Sentiment Classification.
  • 4
    ASPE: Aspect Sentiment Pair Extraction.
  • 5
    CSPE: Category Sentiment Pair Extraction.